Skip to end of metadata
Go to start of metadata

we instrumented V01 today, and it is on the network.

all sensors worked fine in the ops center this morning, but EC150 is not happy on the tower???

will work on servicing problem sensors tomorrow.

No sign of Samortechnica today, tomorrow is a holiday, so I don't expect them here then either.

2 Comments

  1. For whatever reason, the EC150 looks fine at the moment.  All the other sensors on that tower look good too.

     

  2. The only issue is the sample rate for the EC100, its at 10Hz instead of 20.